OpenStudy (anonymous):

Find the rectangular coordinates of (2/3, -2pi/3) So far, I have gotten x=1/3 and y=√3/3 but I do not know if either would be negative because 2pi/3 is in the 2nd quadrant so it would usually be negative, but the negative sign (-) is already in front of it so would it just cancel out? I do not know. Can you help me?

OpenStudy (anonymous):

I got \[ \left\{-\frac{1}{3},-\frac{1}{\sqrt{3 }}\right\} \] Do you need explanation?

OpenStudy (anonymous):

\[ x= \rho \cos(\theta)\\ y= \rho \sin (\theta)\\ \]

OpenStudy (anonymous):

\[ \rho= \frac 2 3\\ \theta=- \frac { 2 \pi}3 \]

OpenStudy (anonymous):

I know all that. So if there is a negative sign in front of a radian that is in quadrant 2, (like 2pi/3 is) then nothing happens because the negative sign is already present? or do the 2 negatives cancel out?

OpenStudy (anonymous):

\[ \cos( -\frac {2 \pi} 3)=\cos(\frac {2 \pi} 3)=-\frac 1 2\\ \sin( -\frac {2 \pi} 3)=-\sin(\frac {2 \pi} 3)=-\frac{\sqrt{3}}{2}\\ \]

OpenStudy (anonymous):

You do not have to worry about signs, just compute correctly the sin and cos as I did above.

OpenStudy (anonymous):

How is your background in Trigonometry?

OpenStudy (anonymous):

This is my second semester in Pre Calculus 2 (and this semester is just trig) but I am on the last unit and actually, the assignment now before the final.

OpenStudy (anonymous):

Did you understand how I computed the sin and the cos above?

OpenStudy (anonymous):

Yes I did. But what I was asking never clarified in the course. And I am in an online class so I have no teacher to ask.

OpenStudy (anonymous):

You have a teacher right now.

OpenStudy (anonymous):

Go to my online test and practice some trig and calc questions: http://mathonline.missouri.edu

OpenStudy (anonymous):

Ok! And thank you so much for the help :)